# On the degrees of freedom of spatially covariant vector field theory

### v1

@@ I Introduction
The purpose of this work is to ask a basic question: is it possible to construct a vector field theory that propagates only the two transverse DOFs [^1] This is also inspired by the recent progress of building gravitational theories with only two DOFs such as [42, 43, 44].? Clearly, enforcing both Lorentz invariance and $\mathrm{U}(1)$ invariance recovers the Maxwell theory for a free vector field. Therefore, to explore possible generalizations, we relax these symmetries and require only spatial covariance, without assuming $\mathrm{U}(1)$ invariance. As a first step, we neglect gravitational effects and work in flat spacetime.
@@ II The spatially covariant vector field theory
In this work, we require that contractions occur only among derivative terms themselves. Terms involving contractions between derivatives and the undifferentiated vector field, such as $A^{i}\partial_{i}A_{0}$, $A^{i}\dot{A}_{i}$ and $A^{i}A^{j}\partial_{i}A_{j}$ will not be considered. This leaves two admissible monomials, $\dot{A}_{0}$ and $\partial_{i}A^{i}$, and the Lagrangian is given by
@@
where again all coefficients are general functions of $A_{0}$ and $\bar{X}$, i.e., $f_{m,n}=f_{m,n}(A_{0},\bar{X})$ and $h_{m,n}=h_{m,n}(A_{0},\bar{X})$. In the second line of (7), the coefficients are denoted as $h_{m,n}$ merely for notational brevity.
@@ III Hamiltonian and primary constraints
For our purpose to find the condition such that the theory propagates only two DOFs, a necessary condition is that the Hessian matrix is degenerate. In order to have $\det\mathcal{H}^{\mu\nu}=0$, we have two possibilities:
@@
Counting number of DOFs can be well performed in the Hamiltonian analysis. With $f_{4,1}=0$, the canonical momenta defined by $\Pi^{\mu}=\delta\mathcal{L}/\delta\dot{A}_{\mu}$ are
$\Pi^{0}$ $=f_{3,1}+f_{4,2}\partial_{i}A^{i},$ (10)
$\Pi^{i}$ $=2h_{4,1}\dot{A}^{i}+h_{4,3}\partial^{i}A_{0}.$ (11)
As expected, $\Pi^{0}$ in (10) contains no time derivatives of the fields, leading to a primary constraint
$\phi_{1}\equiv\Pi^{0}-(f_{4,2}\partial_{i}A^{i}+f_{3,1})\approx 0.$ (12)
Here and throughout this work, $\approx$” represents “weak equality”, which holds on the constraint surface $\Gamma_{p}$ defined by the constraints in phase space.
@@
$\mathcal{H}_{\mathrm{C}}=\Pi^{0}\dot{A}_{0}+\Pi^{i}\dot{A}_{i}-\mathcal{L}.$ (13)
By solving $\dot{A}_{i}$ from (11) and using the primary constraint (12), the canonical Hamiltonian can be expressed as a function of $\Pi^{i}$, $A_{0}$ and $A_{i}$,
$\mathcal{H}_{\mathrm{C}}$ $=\mathcal{K}(\Pi^{i})^{2}+\mathcal{W}_{i}^{(0)}\Pi^{i}+\mathcal{V}^{(0,i)},$ (14)
where $\mathcal{K}$, $\mathcal{W}_{i}^{(0)}$ and $\mathcal{V}^{(0,i)}$ are functions that do not contain $\Pi^{i}$, given by
@@
$F,G]=\int\mathrm{d}^{3}z\left(\frac{\delta F}{\delta A^{I}(\vec{z})}\frac{\delta G}{\delta\Pi_{I}(\vec{z})}-\frac{\delta F}{\delta\Pi_{I}(\vec{z})}\frac{\delta G}{\delta A^{I}(\vec{z})}\right).$ (20)
@@
Note that (21) may either determine the Lagrange multipliers $\{\lambda^{I}\}$ or reduce to relations among the phase space variables $\{A_{\mu},\Pi^{\mu}\}$. In the latter case, there may be secondary constraints if the relations are independent of the primary constraints. Hence, it is necessary to evaluate the Poisson brackets among the primary constraints as well as the Poisson brackets between the primary constraints and the canonical Hamiltonian.
@@
$\mathcal{C}\equiv\dot{\phi}_{1}=\mathcal{A}\partial_{i}\Pi^{i}+\mathcal{B}(\Pi^{i})^{2}+\mathcal{C}_{i}^{(0,i)}\Pi^{i}+\mathcal{D}^{(0,i)},$ (25)
where $(\Pi^{i})^{2}\equiv\Pi_{i}\Pi^{i}$ and $\mathcal{A},\mathcal{B},\mathcal{C}_{i},\mathcal{D}$ are functions involving no $\Pi^{i}$, given by
@@ IV The first degeneracy condition
$\mathcal{F}_{1}=-\frac{\partial\mathcal{A}}{\partial A_{0}}\partial_{i}\Pi^{i}-\frac{\partial\mathcal{B}}{\partial A_{0}}(\Pi^{i})^{2}-\frac{\partial\mathcal{C}_{i}^{(0,i)}}{\partial A_{0}}\Pi^{i}-\frac{\partial\mathcal{D}^{(0,i)}}{\partial A_{0}}+\left(\frac{\partial f_{4,2}}{\partial A_{k}}\partial_{i}A^{i}+\frac{\partial f_{3,1}}{\partial A_{k}}\right)(2\mathcal{B}\Pi_{k}+\mathcal{C}_{k}^{(0,i)}),$ (34)
$\mathcal{F}^{i}=$ $\left[\frac{\partial}{\partial A_{0}}\left(\frac{h_{4,3}}{2h_{4,1}}\right)+\frac{\partial}{\partial A_{0}}\left(\frac{h_{4,3}}{2h_{4,1}}\right)-2\frac{\partial f_{4,2}}{\partial A_{0}}\mathcal{K}+2f_{4,2}\frac{\partial\mathcal{K}}{\partial A_{0}}\right]\Pi^{i}+6\frac{\partial}{\partial A_{0}}\left(h_{4,2}-\frac{(h_{4,3})^{2}}{4h_{4,1}}\right)\partial^{i}A_{0}$
$+\frac{\partial f_{4,2}}{\partial A_{0}}\mathcal{W}^{i(0)}+f_{4,2}\frac{\partial\mathcal{W}^{i(0)}}{\partial A_{0}}+\left(\frac{\partial f_{4,2}}{\partial A_{i}}\partial_{l}A^{l}+\frac{\partial f_{3,1}}{\partial A_{i}}\right)\mathcal{A}+(2\mathcal{B}\Pi^{i}+\mathcal{C}^{i(0,i)})f_{4,2}$
@@
(39) are 3 differential equations for the coefficients, which then constitute the explicit form of the first degeneracy condition. After plugging (11), i.e., $\Pi^{i}=2h_{4,1}\partial_{0}A^{i}+h_{4,3}\partial^{i}A_{0}$, we can solve
@@
where $C_{1}$ and $C_{2}$ are constants. As we have mentioned before, in this work we focus on the case in which the coefficients are polynomials of $A_{0}$ and $\bar{X}$. With this assumption, there is no other possibility than that $h_{4,1}$ is a constant. In other words, we must choose $C_{1}=0$ in the above solutions.
@@
where $\bar{f}_{2}$ and $\bar{f}_{2}$ are general polynomials of $\bar{X}$. For $\mathcal{L}_{3}$, we find that $f_{3,1}$ cannot dependent on $\bar{X}$ while can be a general polynomial of $A_{0}$, while $f_{3,2}$ must be linear in $A_{0}$. Thus we can write
@@
where the coefficients are given as in (41), (42) and (46). As a result, we can rewrite $\dot{\phi}$ as $\dot{\phi}=\mathcal{A}\partial_{i}\Pi^{i}+\mathcal{D}^{(0,i)}$.
@@
where again “XX” denotes generally non-vanishing entries. If no additional constraints appear, (49) implies that all three constraints $\{\phi_{1},\dot{\phi}_{1},\ddot{\phi}_{1}\}$ are second-class, leading formally to 2.5 DOFs. While fractional DOFs can appear in Lorentz-violating theories as what we are discussing, our objective is to identify conditions under which the extra half DOF is fully removed so that only two remain. Thus we need additional conditions in order to reduce the number of DOF to 2.
@@
$\det B^{ab}(\vec{x},\vec{y})=0.$ (50)
We therefore refer to (50) as the second degeneracy condition. It can be shown that $\det B^{ab}$ is proportional to the Poisson brackets $\dot{\phi}_{1},\dot{\phi}_{1}]$ and $\phi_{1},\ddot{\phi}_{1}]$, while has nothing to do with $\dot{\phi}_{1},\ddot{\phi}_{1}]$ or $\ddot{\phi}_{1},\ddot{\phi}_{1}]$. Hence the condition (50) can be realized with either of the following two choices:
@@
Either (51) or (52) is satisfied, we have $\det B^{ab}=0$. We refer to (51) and (52) as the two branches of the second degeneracy condition.
@@ V.1.1 Case 1
$\Phi=\int\mathrm{d}^{3}x\,\phi^{a}(\vec{x})\mathcal{V}_{a}(\vec{x}),$ (66)
which is clear first-class as $\left[\Phi,\phi^{a}(\vec{x})\right]=0$.
For our purpose, we do not need the explicit expression for the new first-class constraint. Instead, we formally denote it as $\tilde{\phi}_{1}$ and the rest two newly combined constraints as $\tilde{\phi}_{2}$ and $\tilde{\phi}_{3}$. Hence the new Dirac matrix becomes
@@ V.1.2 Case 2
Let us now focus on the other case, i.e., when $\mathcal{A}\neq 0$. In this case, the second degeneracy condition (56) implies that quantities in the square brackets if (54) and (55) must be vanishing. This case seems more complicated, especially when explicitly substituting conjugate momenta results in higher order differential equations.
@@
$\mathcal{L}_{3}=f_{3,1}\dot{A}_{0}+\left(\bar{f}_{3,2}\bar{X}+\tilde{f}_{3.2}\right)\partial_{i}A^{i}\simeq\left(\bar{f}_{3,2}\bar{X}+\tilde{f}_{3.2}\right)\partial_{i}A^{i},$ (69)
@@
$h_{4,1}$ $=\alpha_{2},\quad f_{4,2}=\beta_{2}\quad,h_{4,3}=\gamma_{2},$ (71)
@@ V.2 Branch 2
$\ddot{\phi}_{1}$ $=\mathcal{C}_{k}^{*(0,i)}\Pi^{k}+\mathcal{D}^{*(0,i)},$ (76)
@@
$\tilde{\mathcal{F}}_{1}=-\frac{\partial\mathcal{C}^{*k(0,i)}}{\partial A_{0}}\Pi_{k}-\frac{\partial\mathcal{D}^{*(0,i)}}{\partial A_{0}}-\frac{\partial f_{4,2}}{\partial A_{k}}\partial_{i}A^{i}\mathcal{C}_{k}^{*(0,i)}-\frac{\partial f_{3,1}}{\partial A_{k}}\mathcal{C}_{k}^{*(0,i)},$ (80)
@@
By solving (82) with the substitution of $\Pi^{i}=2h_{4,1}\partial_{0}A^{i}+h_{4,3}\partial^{i}A_{0}$, we can get the conditions on the coefficients.
Solving (82), which consists of two differential equations for the coefficients, is generally difficult. Nevertheless, we can discuss two special cases by assuming that the coefficients depend only on either $A_{0}$ or $\bar{X}$ only, which simplifies the analysis. As we will see below, one case yields four second-class constraints, while the other case yields two first-class constraints.

### v2

@@ I Introduction
The purpose of this work is to ask a basic question: is it possible to construct a vector field theory that propagates only the two transverse DOFs [^1] This is also inspired by the recent progress of building gravitational theories with only two DOFs such as [42, 43, 44].? Clearly, enforcing both Lorentz invariance and $\mathrm{U}(1)$ invariance recovers the Maxwell theory for a free vector field. Therefore, to explore possible generalizations, we relax these symmetries and require only spatial covariance, without assuming $\mathrm{U}(1)$ invariance. As a first step, we neglect gravitational effects and work in flat spacetime. This setup can be viewed as a local limit of a foliated cosmological spacetime. Locally the metric is approximately flat on scales much smaller than the curvature scale, while a cosmological evolution can still motivate a preferred time slicing and hence the breaking of time diffeomorphisms. The relevant constraint structure in the vector sector can still be captured in this setup. The analysis in this work therefore serves as a “decoupling limit” of the theory with gravitational effects.
This flat-background analysis should be viewed as a controlled limit of a preferred-foliation cosmological setting: locally (or in the high-frequency/decoupling regime) the background metric can be approximated by Minkowski while the reduced symmetry structure—spatial covariance with broken time diffeomorphisms—still captures the relevant constraint mechanism in the vector sector.
@@ II The spatially covariant vector field theory
In this work, we require that contractions occur only among derivative terms themselves. Terms involving contractions between derivatives and the undifferentiated vector field, such as $A^{i}\partial_{i}A_{0}$, $A^{i}\dot{A}_{i}$ and $A^{i}A^{j}\partial_{i}A_{j}$ will not be considered. This restriction is primarily adopted for tractability and for isolating the degeneracy structure in the simplest nontrivial operator basis. Allowing additional operators (such as $A^{i}\partial_{i}A_{0}$, $A^{i}\dot{A}_{i}$ or $A^{i}A^{j}\partial_{i}A_{j}$) enlarges the basis substantially and in general modifies the momentum dependence of the primary constraint and the subsequent constraint chain, so that additional degeneracy conditions may be required to eliminate the longitudinal mode. The Hamiltonian strategy developed in this work can in principle be extended to the fully general basis, but the systematic classification becomes considerably broader. The present work therefore focuses on the above minimal class to obtain explicit, classifiable two-DOF theories. With this assumption, we are left with two admissible monomials, $\dot{A}_{0}$ and $\partial_{i}A^{i}$, and the Lagrangian is given by
@@
where again all coefficients are general functions of $A_{0}$ and $\bar{X}$, i.e., $f_{m,n}=f_{m,n}(A_{0},\bar{X})$ and $h_{m,n}=h_{m,n}(A_{0},\bar{X})$. In the second line of (7), we denote the corresponding coefficients by $h_{m,n}$ purely for notational convenience. No physical distinction is implied: all $f_{m,n}$ and $h_{m,n}$ are general functions of $A_{0}$ and $\bar{X}$. The split only helps us track the operators involving the $\dot{A}_{i}$-type and $\partial_{i}A_{0}$-type contractions when deriving conjugate momenta and degeneracy conditions.
@@ III Hamiltonian and primary constraints
To find the condition under which the theory propagates only two DOFs, a necessary condition is that the Hessian matrix is degenerate. In order to have $\det\mathcal{H}^{\mu\nu}=0$, we have two possibilities:
@@
Counting number of DOFs can be well performed in the Hamiltonian analysis. With $f_{4,1}=0$, the canonical momenta defined by $\varPi^{\mu}=\delta\mathcal{L}/\delta\dot{A}_{\mu}$ are
$\varPi^{0}$ $=f_{3,1}+f_{4,2}\partial_{i}A^{i},$ (10)
$\varPi^{i}$ $=2h_{4,1}\dot{A}^{i}+h_{4,3}\partial^{i}A_{0}.$ (11)
As expected, $\varPi^{0}$ in (10) contains no time derivatives of the fields, leading to a primary constraint
$\phi_{1}\equiv\varPi^{0}-(f_{4,2}\partial_{i}A^{i}+f_{3,1})\approx 0.$ (12)
Here and throughout this work, $\approx$” represents “weak equality”, which holds on the constraint surface $\varGamma_{\mathrm{p}}$ defined by the constraints in phase space.
@@
$\mathcal{H}_{\mathrm{C}}=\varPi^{0}\dot{A}_{0}+\varPi^{i}\dot{A}_{i}-\mathcal{L}.$ (13)
By solving $\dot{A}_{i}$ from (11) and using the primary constraint (12), the canonical Hamiltonian can be expressed as a function of $\varPi^{i}$, $A_{0}$ and $A_{i}$,
$\mathcal{H}_{\mathrm{C}}$ $=\mathcal{K}(\varPi^{i})^{2}+\mathcal{W}_{i}^{(0)}\varPi^{i}+\mathcal{V}^{(0,i)},$ (14)
where $\mathcal{K}$, $\mathcal{W}_{i}^{(0)}$ and $\mathcal{V}^{(0,i)}$ are functions that do not contain $\varPi^{i}$. Here the superscripts indicate which configuration variables the coefficients depend on. That is, $(0)$ denotes dependence only on $A_{0}$ (and its spatial derivatives), while $(0,i)$ denotes dependence on both $A_{0}$ and $A_{i}$ (and their spatial derivatives). In (14), $\mathcal{K}$, $\mathcal{W}_{i}^{(0)}$ and $\mathcal{V}^{(0,i)}$ are given by
@@
$F,G]=\int\mathrm{d}^{3}z\left(\frac{\delta F}{\delta A^{I}(\vec{z})}\frac{\delta G}{\delta\varPi_{I}(\vec{z})}-\frac{\delta F}{\delta\varPi_{I}(\vec{z})}\frac{\delta G}{\delta A^{I}(\vec{z})}\right).$ (20)
@@
Note that (21) may either determine the Lagrange multipliers $\{\lambda^{I}\}$ or reduce to relations among the phase space variables $\{A_{\mu},\varPi^{\mu}\}$. In the latter case, there may be secondary constraints if the relations are independent of the primary constraints. Hence, it is necessary to evaluate the Poisson brackets among the primary constraints as well as the Poisson brackets between the primary constraints and the canonical Hamiltonian.
@@
$\mathcal{C}\equiv\dot{\phi}_{1}=\mathcal{A}\partial_{i}\varPi^{i}+\mathcal{B}(\varPi^{i})^{2}+\mathcal{C}_{i}^{(0,i)}\varPi^{i}+\mathcal{D}^{(0,i)},$ (25)
where $(\varPi^{i})^{2}\equiv\varPi_{i}\varPi^{i}$ and $\mathcal{A},\mathcal{B},\mathcal{C}_{i},\mathcal{D}$ are functions involving no $\varPi^{i}$, given by
@@ IV The first degeneracy condition
$\mathcal{F}_{1}=-\frac{\partial\mathcal{A}}{\partial A_{0}}\partial_{i}\varPi^{i}-\frac{\partial\mathcal{B}}{\partial A_{0}}(\varPi^{i})^{2}-\frac{\partial\mathcal{C}_{i}^{(0,i)}}{\partial A_{0}}\varPi^{i}-\frac{\partial\mathcal{D}^{(0,i)}}{\partial A_{0}}+\left(\frac{\partial f_{4,2}}{\partial A_{k}}\partial_{i}A^{i}+\frac{\partial f_{3,1}}{\partial A_{k}}\right)(2\mathcal{B}\varPi_{k}+\mathcal{C}_{k}^{(0,i)}),$ (34)
$\mathcal{F}^{i}=$ $\left[\frac{\partial}{\partial A_{0}}\left(\frac{h_{4,3}}{2h_{4,1}}\right)+\frac{\partial}{\partial A_{0}}\left(\frac{h_{4,3}}{2h_{4,1}}\right)-2\frac{\partial f_{4,2}}{\partial A_{0}}\mathcal{K}+2f_{4,2}\frac{\partial\mathcal{K}}{\partial A_{0}}\right]\varPi^{i}+6\frac{\partial}{\partial A_{0}}\left(h_{4,2}-\frac{(h_{4,3})^{2}}{4h_{4,1}}\right)\partial^{i}A_{0}$
$+\frac{\partial f_{4,2}}{\partial A_{0}}\mathcal{W}^{i(0)}+f_{4,2}\frac{\partial\mathcal{W}^{i(0)}}{\partial A_{0}}+\left(\frac{\partial f_{4,2}}{\partial A_{i}}\partial_{l}A^{l}+\frac{\partial f_{3,1}}{\partial A_{i}}\right)\mathcal{A}+(2\mathcal{B}\varPi^{i}+\mathcal{C}^{i(0,i)})f_{4,2}$
@@
(39) are 3 differential equations for the coefficients, which then constitute the explicit form of the first degeneracy condition. After plugging (11), i.e., $\varPi^{i}=2h_{4,1}\partial_{0}A^{i}+h_{4,3}\partial^{i}A_{0}$, we can solve
@@
where $C_{1}$ and $C_{2}$ are constants. As we have mentioned before, in this work we focus on the case in which the coefficients are polynomials of $A_{0}$ and $\bar{X}$. This ansatz is adopted to make the solution space of the (functional) degeneracy conditions explicitly classifiable and to exhibit closed-form families of 2 DOF Lagrangians. The degeneracy conditions themselves can be formulated for more general coefficient functions (e.g. analytic functions admitting local series expansions around backgrounds), and the Hamiltonian/constraint logic below does not rely on polynomiality except when solving the conditions explicitly. With this assumption, there is no other possibility than that $h_{4,1}$ is a constant. In other words, we must choose $C_{1}=0$ in the above solutions.
@@
where $\bar{f}_{2}$ and $\tilde{f}_{2}$ are general polynomials of $\bar{X}$. For $\mathcal{L}_{3}$, we find that $f_{3,1}$ cannot depend on $\bar{X}$ but can be a general polynomial of $A_{0}$, whereas $f_{3,2}$ must be linear in $A_{0}$. Thus we can write
@@
where the coefficients are given as in (41), (42) and (46). As a result, we can rewrite $\dot{\phi}$ as $\dot{\phi}=\mathcal{A}\partial_{i}\varPi^{i}+\mathcal{D}^{(0,i)}$.
@@
where again “XX” denotes generally non-vanishing entries. If no additional constraints appear, (49) implies that all three constraints $\{\phi_{1},\dot{\phi}_{1},\ddot{\phi}_{1}\}$ are second-class, leading formally to 2.5 DOFs. This can be understood as follows. For finite-dimensional systems, an odd-dimensional antisymmetric matrix has vanishing determinant and hence cannot be invertible (nondegenerate). However, this does not directly apply to $B_{ab}(\vec{x},\vec{y})$ in (49). In field theory, the object $B_{ab}(\vec{x},\vec{y})$ in (49) should be regarded as a differential operator acting on test functions, rather than a finite-dimensional matrix. Indeed, in field theories it is possible for an odd number of constraints (per spatial point) to be genuinely second-class when the Poisson brackets contain spatial derivatives of $\delta$-functions and the corresponding operator is invertible after imposing appropriate boundary conditions. A classical example is the Floreanini-Jackiw chiral boson, which is commonly described as carrying one-half physical DOF [48] (see also [49]). Similar phenomena have also been discussed in Lorentz violating field theories, such as Hořava-Lifshitz gravity [50] and spatially covariant gravity theories [51, 44, 52, 53], where an additional condition is required in general to ensure that the would-be half mode is absent and only the desired propagating modes remain. With this understanding, (49) indicates that if $B_{ab}(\vec{x},\vec{y})$ is invertible as an operator on the space of test functions, then $\{\phi_{1},\dot{\phi}_{1},\ddot{\phi}_{1}\}$ are second-class and the theory exhibits a residual half DOF, so that the total number of DOFs is formally 2.5. Our goal is precisely to exclude this possibility and ensure that only the two transverse polarizations propagate.
@@
$\det B^{ab}(\vec{x},\vec{y})=0,$ (50)
which means there exists a nontrivial null vector $V_{b}(\vec{y})$ such that $\int\mathrm{d}^{3}y\,B^{ab}(\vec{x},\vec{y})V_{b}(\vec{y})=0$. We therefore refer to (50) as the second degeneracy condition. It can be shown that $\det B^{ab}$ is proportional to the Poisson brackets $\dot{\phi}_{1},\dot{\phi}_{1}]$ and $\phi_{1},\ddot{\phi}_{1}]$, while it has nothing to do with $\dot{\phi}_{1},\ddot{\phi}_{1}]$ or $\ddot{\phi}_{1},\ddot{\phi}_{1}]$. Hence the condition (50) can be realized with either of the following two choices:
@@
If either (51) or (52) is satisfied, we have $\det B^{ab}=0$. We refer to (51) and (52) as the two branches of the second degeneracy condition.
@@ V.1.1 Case 1
$\varPhi=\int\mathrm{d}^{3}x\,\phi^{a}(\vec{x})\mathcal{V}_{a}(\vec{x}),$ (66)
which is clearly first-class as $\left[\varPhi,\phi^{a}(\vec{x})\right]=0$.
For our purposes, we do not need the explicit expression for the new first-class constraint. Instead, we formally denote it as $\tilde{\phi}_{1}$ and the rest two newly combined constraints as $\tilde{\phi}_{2}$ and $\tilde{\phi}_{3}$. Hence the new Dirac matrix becomes
@@ V.1.2 Case 2
Let us now focus on the other case, i.e., when $\mathcal{A}\neq 0$. In this case, the second degeneracy condition (56) implies that quantities in the square brackets in (54) and (55) must be vanishing. This case seems more complicated, especially when explicitly substituting conjugate momenta results in higher order differential equations.
@@
$\mathcal{L}_{3}=f_{3,1}\dot{A}_{0}+\left(\bar{f}_{3,2}\bar{X}+\tilde{f}_{3.2}\right)\partial_{i}A^{i}\simeq\left(\bar{f}_{3,2}\bar{X}+\tilde{f}_{3,2}\right)\partial_{i}A^{i},$ (69)
@@
$h_{4,1}$ $=\alpha_{2},\quad f_{4,2}=\beta_{2},\quad h_{4,3}=\gamma_{2},$ (71)
@@ V.2 Branch 2
$\ddot{\phi}_{1}$ $=\mathcal{C}_{k}^{*(0,i)}\varPi^{k}+\mathcal{D}^{*(0,i)},$ (76)
@@
$\tilde{\mathcal{F}}_{1}=-\frac{\partial\mathcal{C}^{*k(0,i)}}{\partial A_{0}}\varPi_{k}-\frac{\partial\mathcal{D}^{*(0,i)}}{\partial A_{0}}-\frac{\partial f_{4,2}}{\partial A_{k}}\partial_{i}A^{i}\mathcal{C}_{k}^{*(0,i)}-\frac{\partial f_{3,1}}{\partial A_{k}}\mathcal{C}_{k}^{*(0,i)},$ (80)
@@
By solving (82) with the substitution of $\varPi^{i}=2h_{4,1}\partial_{0}A^{i}+h_{4,3}\partial^{i}A_{0}$, we can get the conditions on the coefficients.
Solving (82), which consists of two differential equations for the coefficients, is generally difficult. Nevertheless, we can discuss two special cases by assuming that the coefficients depend either on $A_{0}$ or $\bar{X}$ only, which simplifies the analysis. As we will see below, one case yields four second-class constraints, while the other case yields two first-class constraints.